OpenStudy (anonymous):

The half life of carbon-14 is 5,730 years. Assume that a typical male has 35 pounds of carbon-14 in his body at death. What is the equation that describes this decay pattern?

OpenStudy (anonymous):

it's an exponential decay. Do you know how to write an exponential decay equation? How about convert a half life into a rate constant?

OpenStudy (anonymous):

\[C(t)=C _{0} e^{-kt}\]

OpenStudy (anonymous):

@dkebler we are learning it without an e, so I was confused when I was trying to look for the equation... only one that I know about it is the y=p(1+r)^x which is the exponential growth equation formula..

OpenStudy (anonymous):

now how is k related to half life???? Half life is really just an understandable way to express a rate constant. It' easy for folks to understand that in 5700 years half will be gone. \[k=\ln2/t_{1/2}\]

OpenStudy (anonymous):

it is not necessary to use base e only more convenient.

OpenStudy (anonymous):

your (1+r) is the base

OpenStudy (anonymous):

r is related to the half life

OpenStudy (anonymous):

\[C(t)=C _{0}(1+r)^t\]

OpenStudy (anonymous):

now you have to figure out r from the half life

OpenStudy (anonymous):

can you do that on your own?

OpenStudy (anonymous):

Oh yeah, that's where I got stuck on... to figure out r..

OpenStudy (anonymous):

and so, do you use C(t)=C0(1+r)t to solve this question?

OpenStudy (anonymous):

ok here is the trick. Half life means time for half to go away

OpenStudy (anonymous):

so you divide 2 by times..?

OpenStudy (anonymous):

\[1/2(C_{0}=C_{0}(1+r){hl}\]

OpenStudy (anonymous):

the Co's cancel you have to take the log of both sides hl is your half life a number you already know then you can solve for r

OpenStudy (anonymous):

y=1/2(5730(1+r)^x...?

OpenStudy (anonymous):

whoops can't solve for r that way, just a moment

OpenStudy (anonymous):

so for example, if I want to figure out how much carbon 14 will be remaining after 1000 years, do I just plug 1000 into x?

OpenStudy (anonymous):

just take the hl root of 1/2 and subtract 1

OpenStudy (anonymous):

first must find r then you can make a prediction

OpenStudy (anonymous):

use the form I gave you X is confusing. t makes more sense. But find r first

OpenStudy (anonymous):

But what do I have to do with the 35 pounds.. well it seems most likely related to figure out what r would be.. but I'm not sure about that

OpenStudy (nincompoop):

a rough estimate

OpenStudy (anonymous):

(1/2)^(1/5730)=1+r

OpenStudy (anonymous):

r=-0.987 ..?

OpenStudy (anonymous):

now you have \[C(t)=35(1+r)^t\]

OpenStudy (anonymous):

y=35(1.99)^t, is that what you are trying to say..?

OpenStudy (anonymous):

hold on should have been 1-r all along as this was a decay (1+r) must be less than 1.

OpenStudy (anonymous):

oh yeah i meant y=35(.01)^t

OpenStudy (anonymous):

just a second.

OpenStudy (anonymous):

ok I get for the base .9998791

OpenStudy (anonymous):

so C(t)=35(.9998791)^t

OpenStudy (anonymous):

now you can put in whatever time you want to making predictions.

OpenStudy (anonymous):

oh I kind of get it now! but then can you explain how did you find the base .99? because I got different answer from what you got;; I think it's because I used different equation, anyways what kind of formula did you use to find out the base .99? this will be my last question!

OpenStudy (anonymous):

the confusion is the 1+r stuff, not needed here the form was just y = p*b^x

OpenStudy (anonymous):

so in this case b = (1/2)^(1/5730)=.9998791

OpenStudy (anonymous):

so after 1000 years later, there will be .00151of carbon-14..?

OpenStudy (anonymous):

no think about it. in 5700 years half is gone, 17.5 remains. In 1000 years it's gonna be like 30 probably

OpenStudy (anonymous):

\[C(10000)=35(.9998791)^(10000)\]

OpenStudy (anonymous):

that's 1000 not 10000 sorry...getting late for me

OpenStudy (anonymous):

oh, I did y=35(.99)^1000. maybe that's why

OpenStudy (anonymous):

yes VERY important to keep as many decimal places in the base as possible cause rounding even a bit will throw all answers off significantly because it is an exponential equation. Ok you good?

OpenStudy (anonymous):

yup! thank you so much for this hard work^^&